# On Robust hypothesis testing with respect to Hellinger distance

### v2

@@ 4 Results: Lower Bound on $\gamma^{*}$
For every $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$, there is a class $\mathcal{P}=\{p_{1},p_{2}\}$ which is not $\gamma$-robustly testable.
We first outline the sketch of the proof. For
@@
    We construct a family of distributions $\mathcal{D}_{1}$ such that for every $p\in\mathcal{D}_{1}$, we have
    $\frac{H^{2}(p,p_{2})}{H^{2}(p,p_{1})}$ $=\frac{\sqrt{2}}{\sqrt{2}-1}.$
    We construct another family of distributions $\mathcal{D}_{2}$ such that for every $p\in\mathcal{D}_{2}$, we have
    $\frac{H^{2}(p,p_{1})}{H^{2}(p,p_{2})}=\frac{\sqrt{2}}{\sqrt{2}-1}.$
@@
    Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from $p$. We show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{2}^{m})\leq\frac{1}{3}$.
@@
    We use Le Cam’s argument to show that the probability of error in distinguishing between $\mathcal{D}_{1}^{m}$ and $\mathcal{D}_{2}^{m}$ is at least $\frac{1}{3}$.
@@
    The above three points imply that if we had a $\gamma$-robust test with $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$ then we could use it to distinguish between $\mathcal{D}_{1}^{m}$ and $\mathcal{D}_{2}^{m}$ with probability of error at most $\frac{1}{3}$ which contradicts point (3). Hence, such a test cannot exist.
@@
Figure 1: $\mathcal{D}_{1}$ is a family of distributions such that all its members are $\frac{\sqrt{2}}{\sqrt{2}-1}$ times farther to $p_{2}$ than to $p_{1}$ in Hellinger distance. Likewise, all the members of $\mathcal{D}_{2}$ are $\frac{\sqrt{2}}{\sqrt{2}-1}$ times farther to $p_{1}$ than to $p_{2}$ in Hellinger distance.
@@
We first construct the families $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ described in the proof sketch. Let $0<b\leq 1$. We will set the value of $b$ later. Define distribution $p_{1}$ as follows.
$p_{1}(x)=\begin{cases}1-b&x\leq 0.5\\ 1+b&0.5<x\leq 1.\end{cases}$
Define distribution $p_{2}$ as follows.
$p_{2}(x)=\begin{cases}1+b&x\leq 0.5\\ 1-b&0.5<x\leq 1\end{cases}$
Divide $[0,1]$ interval into $2N_{m}$ bins of equal size. Let $I_{k}=\left[\frac{k-1}{2N_{m}},\frac{k}{2N_{m}}\right)$ be the $k^{\text{th}}$ bin. Let $a_{1}\geq b$ and $b\leq a_{2}\leq 1+b$. We select a subset $R_{1}\subseteq[N_{m}]$ of size $\left({\frac{b}{a_{1}}}\right)N_{m}$. We also select a subset $R_{2}\subseteq[N_{m}]$ of size $\left(\frac{b}{a_{2}}\right)N_{m}$. Consider the distribution $p^{R_{1},R_{2}}$ obtained by perturbing $p_{1}$.
@@
Let $\mathcal{D}_{1}=\left\{p^{R_{1},R_{2}}:R_{1},R_{2}\subseteq[N_{m}],|R_{1}|=\left({\frac{b}{a_{1}}}\right)N_{m},|R_{2}|=\left({\frac{b}{a_{2}}}\right)N_{m}\right\}$ be the family of distributions (parameterized by the sets $R_{1},R_{2}$) obtained by perturbing $p_{1}$. Now, consider the distribution $\bar{p}^{R_{1},R_{2}}$ obtained by perturbing $p_{2}$.
@@
Let $\mathcal{D}_{2}=\left\{\bar{p}^{R_{1},R_{2}}:R_{1},R_{2}\subseteq[N_{m}],|R_{1}|=\left({\frac{b}{a_{1}}}\right)N_{m},|R_{2}|=\left({\frac{b}{a_{2}}}\right)N_{m}\right\}$ be the family of distributions (parameterized by the sets $R_{1},R_{2}$) obtained by perturbing $p_{2}$.
@@
Figure 2: An example of $p^{R_{1},R_{2}}$ (perturbed around $p_{1}$) and $\bar{p}^{R_{1},R_{2}}$ (perturbed around $p_{2}$) when $R_{1}=\{2,4\}$ and $R_{2}=\{1,3\}$.
We now set $b=1,a_{2}=1$ and let $a_{1}\rightarrow\infty$. For any $p^{R_{1},R_{2}}\in\mathcal{D}_{1}$, its Hellinger distance from $p_{1}$ is given by
$H^{2}(p,p_{1})$ $=\frac{1}{2}\left[(\sqrt{a_{1}}-0)^{2}\left(\frac{1}{2N_{m}}\right)\left(\frac{1}{a_{1}}N_{m}\right)+(\sqrt{2}-1)^{2}\left(\frac{1}{2N_{m}}\right)N_{m}\right]$
@@
Also, its Hellinger distance from $p_{2}$ is given by
$H^{2}(p,p_{2})$ $=\frac{1}{2}\left[(\sqrt{2}-0)^{2}\frac{1}{2N_{m}}\left(1-\frac{1}{a_{1}}\right)N_{m}+(\sqrt{a_{1}}-\sqrt{2})^{2}\frac{1}{2N_{m}}\frac{1}{a_{1}}N_{m}+(\sqrt{1}-0)^{2}\frac{1}{2N_{m}}N_{m}\right]$
@@
$\lim_{a_{1}\rightarrow\infty}\frac{H^{2}(p,p_{2})}{H^{2}(p,p_{1})}$ $=\frac{2+1+1}{4-2\sqrt{2}}$
@@
Likewise for any $\bar{p}^{R_{1},R_{2}}\in\mathcal{D}_{2}$, we have
$\lim_{a_{1}\rightarrow\infty}\frac{H^{2}(p,p_{1})}{H^{2}(p,p_{2})}$ $=\frac{\sqrt{2}}{\sqrt{2}-1}.$
Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from $p$. We now show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{1}^{m})$ is small for an appropriate choice of $N_{m}$. We use [7, Lemma 24], which states that two distributions are close (in TV) if there exists an event $E$ such that both distributions put a large mass on $E$ and conditioned on this event the two distributions are close (in TV). We reproduce the proof for completeness.
Lemma 1 ([7, Lemma 24]).
@@
We use the above lemma with $u=\mathcal{D}_{1}^{m}$, $v=\mathcal{D}_{2}^{m}$ and $E$ the event where every interval $I_{j}$ contains at most one sample. Observe that the event $E$ is invariant under any permutation of $I_{j}$’s. Furthermore if $p$ is picked uniformly at random from the family $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$ then $\mathbb{E}[p]=\text{unif}[0,1]$, i.e. the uniform distribution over $[0,1]$. Using [7, Claim 26], we have ${\mathcal{D}_{1}^{m}}_{|E}={\mathcal{D}_{2}^{m}}_{|E}$. Thus,
$TV(\mathcal{D}_{1}^{m},\mathcal{D}_{1}^{m})\leq 2\mathcal{D}_{1}^{m}(E^{c})+2\mathcal{D}_{2}^{m}(E^{c}).$
@@
$\mathcal{D}_{1}^{m}(E^{c})$ $\geq\left(1-\frac{1}{2N_{m}}\right)\left(1-\frac{2}{2N_{m}}\right)\cdots\left(1-\frac{m-1}{2N_{m}}\right)$
@@
$\approx e^{\frac{-(m-1)^{2}}{N_{m}}}$
If we pick $N_{m}=C(m-1)^{2}$ then $\mathcal{D}_{1}^{m}(E^{c})\geq\frac{11}{12}$ for some appropriate constant $C$. Likewise, we have $\mathcal{D}_{2}^{m}(E^{c})\geq\frac{11}{12}$. Thus, we have $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{2}^{m})\leq\frac{1}{3}$. We now use Le Cam’s Lemma [10, Lemma 1].
@@
We again use this lemma with $u=\mathcal{D}_{1}^{m}$, $v=\mathcal{D}_{2}^{m}$. Thus, for any test $A\subset\mathcal{X}^{m}$ we have
$\max(\max_{p\in\mathcal{D}_{1}}p(A^{c}),\max_{\bar{p}\in\mathcal{D}_{2}}\bar{p}(A))$ $\geq\max(\mathcal{D}_{1}^{m}(A^{c}),\mathcal{D}_{2}^{m}(A))$
$\geq\frac{1-TV(\mathcal{D}_{1}^{m},\mathcal{D}_{2}^{m})}{2}$
@@
Thus, if a $\gamma$-robust test with $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$ existed then we could use it to distinguish between the case where $p\in\mathcal{D}_{1}$ and where $p\in\mathcal{D}_{2}$ with probability of error at most $\frac{1}{3}$. This leads to a contradiction and completes the proof. ∎
@@ 4-A When is the lower bound tight?
Let $\text{supp}(p)$ denote the support of distribution $p$, i.e $\text{supp}(p)=\{x\in\mathcal{X}:p(x)>0\}$. We now show that our lower bound is tight if $\text{supp}(p_{1})\cap\text{supp}(p_{2})=\emptyset$. Let $S_{1}:=\text{supp}(p_{1})$, $S_{2}:=\text{supp}(p_{2})$ and $S:=\mathcal{X}\backslash(S_{1}\cup S_{2})$. In this scenario, any observed sample that is not in either $S_{1}$ or $S_{2}$ is not informative for the problem at hand. Thus, a natural test would be to declare $H_{0}$ (resp. $H_{1}$) if the fraction of samples falling in $S_{1}$ (resp. $S_{2}$) is greater than $\frac{1}{2}$. Let
$T(X^{n})=\frac{1}{n}\sum_{i=1}^{n}\mathrm{1}[X_{i}\in S_{1}]-\mathrm{1}[X_{i}\in S_{2}]$
@@
$\mathbb{E}[T(X^{n})]=p(S_{1})-p(S_{2})$
We would like to show that if $\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{1})\leq H^{2}(p,p_{2})$ then $\mathbb{E}[T]>0$ and if $H^{2}(p,p_{1})\leq\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{2})$ then $\mathbb{E}[T]<0$. We have
$=\frac{1}{2}\sum_{x\in S_{1}}(\sqrt{p(x)}-\sqrt{p_{1}(x)})^{2}+\sum_{x\in S_{2}}(\sqrt{p(x)}-\sqrt{p_{1}(x)})^{2}+\sum_{x\in S}(\sqrt{p(x)}-\sqrt{p_{1}(x)})^{2}$
$\overset{(a)}{=}\frac{1}{2}\left[\sum_{x\in S_{1}}p_{1}(x)-\sum_{x\in S_{1}}2\sqrt{p(x)}\sqrt{p_{1}(x)}+\sum_{x\in S_{1}}p(x)+\sum_{x\in S_{2}}p(x)+\sum_{x\in S}p(x)\right]$
$=\frac{1}{2}\left[2-2\sum_{x\in S_{1}}\left(\sqrt{p(x)}\sqrt{p_{1}(x)}\right)\right]$
@@
where $(a)$ follows from the definition of $S_{1}$, $(b)$ follows from the fact that $\sqrt{p_{1}(x)}\leq 1$ and $(c)$ follows from the concavity of the square root function. Also,
$H^{2}(p,p_{2})$ $=\frac{1}{2}\sum_{x\in\mathcal{X}}(\sqrt{p(x)}-\sqrt{p_{2}(x)})^{2}$
$\leq\frac{1}{2}\sum_{x\in\mathcal{X}}p(x)+p_{2}(x)-2\sqrt{p(x)}\sqrt{p_{2}(x)}$
$=\frac{1}{2}\sum_{x\in\mathcal{X}}p(x)+p_{2}(x)$
@@
$\frac{\sqrt{2}}{\sqrt{2}-1}\left(1-\sqrt{p(S_{1})}\right)\leq\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{1})\leq H^{2}(p,p_{2})\leq 1$
Rearranging terms, we get $p(S_{1})\geq\frac{1}{2}$ or $\mathbb{E}[T(X^{n})]\geq 0$. Similarly, we can show that $\mathbb{E}[T(X^{n})]\leq 0$ under $H_{1}$.
Under the constraint $\text{supp}(p_{1})\cap\text{supp}(p_{2})=\emptyset$, the optimal slack factor for TV distance goes down from 3 to 2. This indicates that our lower bound might not be tight.
5 Some results on testing with respect to symmetric $\chi^{2}$ distance
The symmetric $\chi^{2}$ distance between two distributions $p_{1}$ and $p_{2}$ is defined by
$\chi^{2}(p_{1},p_{2})$ $:=\bigg\|\frac{p_{1}-p_{2}}{\sqrt{p_{1}+p_{2}}}\bigg\|_{2}^{2}.$

### v3

@@ 4 Lower Bound on $\gamma^{*}$
For every $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$, there is a class $\mathcal{P}=\{p_{0},p_{1}\}$ which is not $\gamma$-robustly testable.
We first outline the sketch of the proof. We first fix $p_{0}$ and $p_{1}$. Let $m$ be the number of i.i.d. samples drawn from the true distribution $p$ . For any fixed $m$, we have the following.
@@
    We construct a family of distributions $\mathcal{D}_{0}$ such that for every $p\in\mathcal{D}_{0}$ and any arbitrarily small $\epsilon$, we have
    $\frac{\sqrt{2}}{\sqrt{2}-1}-\epsilon<\frac{H^{2}(p,p_{1})}{H^{2}(p,p_{0})}<\frac{\sqrt{2}}{\sqrt{2}-1}.$
    We construct another family of distributions $\mathcal{D}_{1}$ such that for every $p\in\mathcal{D}_{1}$ and any arbitrarily small $\epsilon$, we have
    $\frac{\sqrt{2}}{\sqrt{2}-1}-\epsilon<\frac{H^{2}(p,p_{0})}{H^{2}(p,p_{1})}<\frac{\sqrt{2}}{\sqrt{2}-1}.$
@@
    Let $\mathcal{D}_{0}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{0}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{1}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. We show that $TV(\mathcal{D}_{0}^{m},\mathcal{D}_{1}^{m})\leq\frac{1}{3}$.
@@
    We use Le Cam’s argument to show that the probability of error in distinguishing between $\mathcal{D}_{0}^{m}$ and $\mathcal{D}_{1}^{m}$ is at least $\frac{1}{3}$.
@@
    The above three points imply that if we had a $\gamma$-robust test with $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$ then we could use it to distinguish between $\mathcal{D}_{0}^{m}$ and $\mathcal{D}_{1}^{m}$ with probability of error at most $\frac{1}{3}$ which contradicts point (3). Hence, such a test cannot exist.
@@
Figure 1: $\mathcal{D}_{0}$ is a family of distributions such that all its members are $\frac{\sqrt{2}}{\sqrt{2}-1}$ times farther to $p_{1}$ than to $p_{0}$ in Hellinger distance. Likewise, all the members of $\mathcal{D}_{1}$ are $\frac{\sqrt{2}}{\sqrt{2}-1}$ times farther to $p_{0}$ than to $p_{1}$ in Hellinger distance.
@@
We first construct the families $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ described in the proof sketch. Let $0<b\leq 1$. We will set the value of $b$ later. Define distribution $p_{0}$ as follows.
$p_{0}(x)=\begin{cases}1-b&x\leq 0.5\\ 1+b&0.5<x\leq 1.\end{cases}$
Define distribution $p_{1}$ as follows.
$p_{1}(x)=\begin{cases}1+b&x\leq 0.5\\ 1-b&0.5<x\leq 1\end{cases}$
Divide $[0,1]$ interval into $2N_{m}$ bins of equal size. Let $I_{k}=\left[\frac{k-1}{2N_{m}},\frac{k}{2N_{m}}\right)$ be the $k^{\text{th}}$ bin. Let $a_{1}\geq b$ and $b\leq a_{2}\leq 1+b$. We select a subset $R_{1}\subseteq[N_{m}]$ of size $\left({\frac{b}{a_{1}}}\right)N_{m}$. We also select a subset $R_{2}\subseteq[N_{m}]$ of size $\left(\frac{b}{a_{2}}\right)N_{m}$. Consider the distribution $p^{R_{1},R_{2}}$ obtained by perturbing $p_{0}$.
@@
Let $\mathcal{D}_{0}=\left\{p^{R_{1},R_{2}}:R_{1},R_{2}\subseteq[N_{m}],|R_{1}|=\left({\frac{b}{a_{1}}}\right)N_{m},|R_{2}|=\left({\frac{b}{a_{2}}}\right)N_{m}\right\}$ be the family of distributions (parameterized by the sets $R_{1},R_{2}$) obtained by perturbing $p_{0}$. Now, consider the distribution $\bar{p}^{R_{1},R_{2}}$ obtained by perturbing $p_{1}$.
@@
Let $\mathcal{D}_{1}=\left\{\bar{p}^{R_{1},R_{2}}:R_{1},R_{2}\subseteq[N_{m}],|R_{1}|=\left({\frac{b}{a_{1}}}\right)N_{m},|R_{2}|=\left({\frac{b}{a_{2}}}\right)N_{m}\right\}$ be the family of distributions (parameterized by the sets $R_{1},R_{2}$) obtained by perturbing $p_{1}$.
@@
Figure 2: An example of $p^{R_{1},R_{2}}$ (perturbed around $p_{0}$) and $\bar{p}^{R_{1},R_{2}}$ (perturbed around $p_{1}$) when $R_{1}=\{2,4\}$ and $R_{2}=\{1,3\}$.
We now set $b=1,a_{2}=1$ and let $a_{1}\rightarrow\infty$. For any $p^{R_{1},R_{2}}\in\mathcal{D}_{0}$, its Hellinger distance from $p_{0}$ is given by
$H^{2}(p,p_{0})$ $=\frac{1}{2}\left[(\sqrt{a_{1}}-0)^{2}\left(\frac{1}{2N_{m}}\right)\left(\frac{1}{a_{1}}N_{m}\right)+(\sqrt{2}-1)^{2}\left(\frac{1}{2N_{m}}\right)N_{m}\right]$
@@
Also, its Hellinger distance from $p_{1}$ is given by
$H^{2}(p,p_{1})$ $=\frac{1}{2}\left[(\sqrt{2}-0)^{2}\frac{1}{2N_{m}}\left(1-\frac{1}{a_{1}}\right)N_{m}+(\sqrt{a_{1}}-\sqrt{2})^{2}\frac{1}{2N_{m}}\frac{1}{a_{1}}N_{m}+(\sqrt{1}-0)^{2}\frac{1}{2N_{m}}N_{m}\right]$
@@
$\lim_{a_{1}\rightarrow\infty}\frac{H^{2}(p,p_{1})}{H^{2}(p,p_{0})}$ $=\frac{2+1+1}{4-2\sqrt{2}}$
@@
Likewise for any $\bar{p}^{R_{1},R_{2}}\in\mathcal{D}_{1}$, we have
$\lim_{a_{1}\rightarrow\infty}\frac{H^{2}(p,p_{0})}{H^{2}(p,p_{1})}$ $=\frac{\sqrt{2}}{\sqrt{2}-1}.$
Let $\mathcal{D}_{0}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{0}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{1}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. We now show that $TV(\mathcal{D}_{0}^{m},\mathcal{D}_{0}^{m})$ is small for an appropriate choice of $N_{m}$. We use [2, Lemma 24], which states that two distributions are close (in TV) if there exists an event $E$ such that both distributions put a large mass on $E$ and conditioned on this event the two distributions are close (in TV). We reproduce the proof for completeness.
Lemma 1 ([2, Lemma 24]).
@@
We use the above lemma with $u=\mathcal{D}_{0}^{m}$, $v=\mathcal{D}_{1}^{m}$ and $E$ the event where every interval $I_{j}$ contains at most one sample. Observe that the event $E$ is invariant under any permutation of $I_{j}$’s. Furthermore, if $p$ is picked uniformly at random from the family $\mathcal{D}_{0}$ or $\mathcal{D}_{1}$ then $\mathbb{E}[p]=\text{unif}[0,1]$, i.e. the uniform distribution over $[0,1]$. Using [2, Claim 26], we have ${\mathcal{D}_{0}^{m}}_{|E}={\mathcal{D}_{1}^{m}}_{|E}$. Thus,
$TV(\mathcal{D}_{0}^{m},\mathcal{D}_{0}^{m})\leq 2\mathcal{D}_{0}^{m}(E^{c})+2\mathcal{D}_{1}^{m}(E^{c}).$
@@
$\mathcal{D}_{0}^{m}(E)$ $\geq\left(1-\frac{1}{2N_{m}}\right)\left(1-\frac{2}{2N_{m}}\right)\cdots\left(1-\frac{m-1}{2N_{m}}\right)$
@@
$\approx e^{\frac{-(m-1)^{2}}{2N_{m}}}$
If we pick $N_{m}=C(m-1)^{2}$ then $\mathcal{D}_{0}^{m}(E)\geq\frac{11}{12}$ for some appropriate constant $C$. Likewise, we have $\mathcal{D}_{1}^{m}(E)\geq\frac{11}{12}$. Thus, we have $TV(\mathcal{D}_{0}^{m},\mathcal{D}_{1}^{m})\leq\frac{1}{3}$. We now use Le Cam’s Lemma [10, Lemma 1].
@@
We again use this lemma with $u=\mathcal{D}_{0}^{m}$, $v=\mathcal{D}_{1}^{m}$. Thus, for any test $A\subset\mathcal{X}^{m}$ we have
$\max(\max_{p\in\mathcal{D}_{0}}p(A^{c}),\max_{\bar{p}\in\mathcal{D}_{1}}\bar{p}(A))$ $\geq\max(\mathcal{D}_{0}^{m}(A^{c}),\mathcal{D}_{1}^{m}(A))$
$\geq\frac{1-TV(\mathcal{D}_{0}^{m},\mathcal{D}_{1}^{m})}{2}$
@@
Thus, if a $\gamma$-robust test with $\gamma<\frac{\sqrt{2}}{\sqrt{2}-1}$ existed then we could use it to distinguish between the case where $p\in\mathcal{D}_{0}$ and where $p\in\mathcal{D}_{1}$ with probability of error at most $\frac{1}{3}$. This leads to a contradiction and completes the proof. ∎
@@ 4-A When is the lower bound tight?
Let $\text{supp}(p)$ denote the support of distribution $p$, i.e $\text{supp}(p)=\{x\in\mathcal{X}:p(x)>0\}$. We now show that our lower bound is tight if $\text{supp}(p_{0})\cap\text{supp}(p_{1})=\emptyset$. Let $S_{1}:=\text{supp}(p_{0})$, $S_{2}:=\text{supp}(p_{1})$ and $S:=\mathcal{X}\backslash(S_{1}\cup S_{2})$. In this scenario, any observed sample that is not in either $S_{1}$ or $S_{2}$ is not informative for the problem at hand. Thus, a natural test would be to declare $H_{0}$ (resp. $H_{1}$) if the fraction of samples falling in $S_{1}$ (resp. $S_{2}$) is greater than $\frac{1}{2}$. Let
$T(X^{m})=\frac{1}{m}\sum_{i=1}^{m}\mathrm{1}[X_{i}\in S_{1}]-\mathrm{1}[X_{i}\in S_{2}]$
@@
$\mathbb{E}[T(X^{m})]=p(S_{1})-p(S_{2})$
We would like to show that if $\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{0})\leq H^{2}(p,p_{1})$ then $\mathbb{E}[T]>0$ and if $H^{2}(p,p_{0})\leq\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{1})$ then $\mathbb{E}[T]<0$. We have
$H^{2}(p,p_{0})$ $=\frac{1}{2}\sum_{x\in\mathcal{X}}(\sqrt{p(x)}-\sqrt{p_{0}(x)})^{2}$
$=\frac{1}{2}\sum_{x\in S_{1}}(\sqrt{p(x)}-\sqrt{p_{0}(x)})^{2}+\sum_{x\in S_{2}}(\sqrt{p(x)}-\sqrt{p_{0}(x)})^{2}+\sum_{x\in S}(\sqrt{p(x)}-\sqrt{p_{0}(x)})^{2}$
$\overset{(a)}{=}\frac{1}{2}\left[\sum_{x\in S_{1}}p_{0}(x)-\sum_{x\in S_{1}}2\sqrt{p(x)}\sqrt{p_{0}(x)}+\sum_{x\in S_{1}}p(x)+\sum_{x\in S_{2}}p(x)+\sum_{x\in S}p(x)\right]$
$=\frac{1}{2}\left[2-2\sum_{x\in S_{1}}\left(\sqrt{p(x)}\sqrt{p_{0}(x)}\right)\right]$
@@
where $(a)$ follows from the definition of $S_{1}$, $(b)$ follows from the fact that $\sqrt{p_{0}(x)}\leq 1$ and $(c)$ follows from the concavity of the square root function. Also,
$\leq\frac{1}{2}\sum_{x\in\mathcal{X}}p(x)+p_{1}(x)-2\sqrt{p(x)}\sqrt{p_{1}(x)}$
$=\frac{1}{2}\sum_{x\in\mathcal{X}}p(x)+p_{1}(x)$
@@
$\frac{\sqrt{2}}{\sqrt{2}-1}\left(1-\sqrt{p(S_{1})}\right)\leq\frac{\sqrt{2}}{\sqrt{2}-1}H^{2}(p,p_{0})\leq H^{2}(p,p_{1})\leq 1$
Rearranging terms, we get $p(S_{1})\geq\frac{1}{2}$ or $\mathbb{E}[T(X^{m})]\geq 0$. Similarly, we can show that $\mathbb{E}[T(X^{m})]\leq 0$ under $H_{1}$.
Under the constraint $\text{supp}(p_{0})\cap\text{supp}(p_{1})=\emptyset$, the optimal slack factor for TV distance goes down from 3 to 2. This indicates that our lower bound might not be tight.
5 Lower bound on testing with respect to symmetric $\chi^{2}$ distance
The symmetric $\chi^{2}$ distance between two distributions $p_{0}$ and $p_{1}$ is defined by
$\chi^{2}(p_{0},p_{1})$ $:=\bigg\|\frac{p_{0}-p_{1}}{\sqrt{p_{0}+p_{1}}}\bigg\|_{2}^{2}.$